OpenStudy (anonymous):

How can I make an equation in slope-intercept form for the line passing through (5, 1) and (–4, 7) ?

OpenStudy (ccswims):

First find the slope (or average rate of change)

OpenStudy (anonymous):

-2/3 ? @ccswims

OpenStudy (ccswims):

switch it, -3/2

OpenStudy (anonymous):

Are you sure?

OpenStudy (ccswims):

-4 - 5/ 7 - 1 = -9/6 = -3/2

OpenStudy (ccswims):

now put it in point-slope form (y - y1) = m(x - x1)

OpenStudy (anonymous):

Oh! Hm.

OpenStudy (ccswims):

Do you need help with that?

OpenStudy (anonymous):

(y - 1) = m (x - 5)

OpenStudy (anonymous):

@ccswims ^

OpenStudy (ccswims):

Remember, our 'm' is our slope So plug in the slope we got earlier for 'm'

OpenStudy (anonymous):

Oh, right! (y - 1) = -3/2 (x - 5)

OpenStudy (ccswims):

good, now distribute the -3/2 to x and -5

OpenStudy (anonymous):

(y - 1) = -3/2x - (-3/2) * 5

OpenStudy (anonymous):

Like that?

OpenStudy (ccswims):

(y - 1) = -3/2x + 7.5

OpenStudy (ccswims):

now add 1 to both sides

OpenStudy (anonymous):

Okay, so what would be the next step?

OpenStudy (anonymous):

Oh, okay!

OpenStudy (anonymous):

y - 1 + 1 = (-3/2)x + 8.5

OpenStudy (ccswims):

good, and -1 + 1 = 0 so... y = -3/2x + 8.5

OpenStudy (anonymous):

y = -3/2 + 8.5

OpenStudy (ccswims):

yup! And now it's in point-slope form :)

OpenStudy (anonymous):

Thanks! But the only problem is that the answer choices all say -2/3 :/ @ccswims

OpenStudy (ccswims):

OH, nevermind, my mistake, it is -2/3

OpenStudy (ccswims):

I did x2 - x1/y2 - y1 when it's supposed to be y2 - y1/x2 - x1

OpenStudy (anonymous):

The answer choices all say 13/3 not 8.5 :( Where did we go wrong?

OpenStudy (ccswims):

yea, instead of doing -3/2 and multiplying it to -5 it would be -2/3 * -5, the answer is

OpenStudy (anonymous):

13/3 ?

OpenStudy (ccswims):

It's A

OpenStudy (anonymous):

I knew it! Well, thanks for your explanation and the help! :)